Suppose that $2000 is invested at a rate of 4.4% , compounded quarterly. Assuming that no withdrawals are made, find the total amount after
7years.Do not round any intermediate computations, and round your answer to the nearest cent.

Answer by rfer(16322)

FV=2000(1+0.044/4)^(4*7)
FV=2000(1.011)^28
FV=2000*1.35841465
FV=2716.83